# Pragmatic Curiosity: A Hybrid Learning-Optimization Paradigm via Active Inference

### v1

@@ Abstract
Many engineering and scientific workflows depend on expensive black-box evaluations, requiring decision-making that simultaneously improves performance and reduces uncertainty. Bayesian optimization (BO) and Bayesian experimental design (BED) offer powerful yet largely separate treatments of goal-seeking and information-seeking, providing limited guidance for hybrid settings where learning and optimization are intrinsically coupled. We propose pragmatic curiosity, a hybrid learning-optimization paradigm derived from active inference, in which actions are selected by minimizing the expected free energy—a single objective that couples pragmatic utility with epistemic information gain. We demonstrate the practical effectiveness and flexibility of pragmatic curiosity on various real-world hybrid tasks, including constrained system identification, targeted active search, and composite optimization with unknown preferences. Across these benchmarks, pragmatic curiosity consistently outperforms strong BO-type and BED-type baselines, delivering higher estimation accuracy, better critical-region coverage, and improved final solution quality.
Active inference, Hybrid learning-optimization
@@ 1 Introduction
Engineering and scientific applications often rely on expensive black-box evaluations to identify high-performing designs or desirable system states. Bayesian optimization (BO) accelerates this process when the primary objective is to reach a specified goal (Shahriari et al., 2016; Frazier, 2018), whereas Bayesian experimental design (BED) prioritizes information acquisition about unknown system parameters (Rainforth et al., 2023). Both methodologies leverage probabilistic models and acquisition criteria that quantify the utility of evaluating unknown configurations, tailored to either optimization or learning objectives. Despite their individual successes and the explosion of research in each field, their disconnection creates a vacuum for a broad class of hybrid problems that routinely require simultaneously both seeking knowledge and achieving goals.
For many real-world applications, such as goal-directed planning (Lookman et al., 2019), environmental monitoring (Konakovic Lukovic et al., 2020), and targeted material design (Matsumoto et al., 2025), learning and optimization are not separate phases but deeply intertwined objectives. This challenge fundamentally arises across tasks with increasing complexity in terms of both epistemic consideration (i.e., from parametric to non-parametric models) and pragmatic evaluation (i.e., from known to unknown goals):
(1) Constrained System Identification, where the epistemic desire of precisely learning a system’s parameters is governed by the pragmatic need to keep experiments within safe or valid operational bounds (e.g., avoiding sensor saturation, or dangerous chemical reaction). This type of task can be found in numerous applications, including environmental monitoring (Konakovic Lukovic et al., 2020) and catalyst design (Zhong et al., 2020).
(2) Targeted Active Search, where the pragmatic objective to discover regions that meet certain criteria (e.g., system failure modes, or specific performance ranges) requires epistemic curiosity to explore the region’s shape, size, and boundaries. Example applications can be found in failure discovery (Ramanagopal et al., 2018) and medical monitoring (Malkomes et al., 2021).
(3) Composite Bayesian Optimization, where the pragmatic goal is to find an optimal design based on a user’s hidden preferences—a task that is impossible without first being epistemically curious about the user’s objectives themselves. Such scenarios commonly arise in simulation-based design (González and Zavala, 2025; Coelho et al., 2025) and A/B testing (Bakshy et al., 2018)
@@
Classically, to address these hybrid problems, practitioners have been forced to choose between specialized tools, and accommodate problem-specific adaptation by leveraging information-gain criteria to enhance optimization and vice versa. On the BO side, Russo and Van Roy (2018) proposed information-directed sampling (IDS) to online optimization problems. Hvarfner et al. (2023) introduced a statistical distance-based active learning (SAL) criterion into the BO loop to actively learn the model hyperparameter even as it searches for the optimum. On the BED (also known as Bayesian active learning, BAL) side, Smith et al. (2023) proposed an expected predictive information gain (EPIG) criterion that focuses on information gain in model predictions, mitigating classical BAL’s tendency to select out-of-distribution or low-relevance queries by accounting for an input data distribution. These methods highlight growing synergy between optimization and learning, but they remain task-specific, and rarely generalize across categories.
In this paper, we propose pragmatic curiosity: a hybrid learning-optimization paradigm derived from active inference (AIF) (Friston, 2010; Friston et al., 2017). AIF prescribes action selection by minimizing expected free energy (EFE), a single objective that combines (i) a pragmatic term that favors preferred outcomes and (ii) an epistemic term that favors information gain. We prove that EFE minimization provides a unified perspective of various acquisition strategies: by specifying preferences, observation models, and approximations, the resulting criterion recovers BO-like and BED-like behaviors as limiting regimes.
Under this paradigm, seeking knowledge and achieving goals are not treated as competing objectives, but as two inseparable facets of a single imperative to minimize EFE. These two drives are balanced by a coefficient called curiosity, which sets the trade-off between learning and optimization. The formal role of curiosity in guaranteeing both self-consistent learning (i.e., posterior convergences to the truth) and no-regret optimization (i.e., with bounded cumulative regret) is provided theoretically in Li et al. (2026).
This paper, instead, demonstrates the practical effectiveness and flexibility of this paradigm in handling a broad class of complex hybrid problems often ignored by standard methods, including tasks with evolving goals (where conditions change over time) and implicit goals (where objectives are not defined a priori). We conduct experiments structured around the three problem classes mentioned above, drawing on applications that include environmental monitoring in plume fields (Konakovic Lukovic et al., 2020), failure detection in autonomous driving scenarios (Ramanagopal et al., 2018), and distributed energy resource allocation in power grids (Kianmehr et al., 2019).
The empirical results reveal a consistent pattern of superior performance, demonstrating our framework’s advantages in its ability to solve complex, hybrid objectives. In constrained system identification tasks, our algorithm achieved near-perfect estimation accuracy while requiring up to 40% fewer queries than other methods. For targeted active search tasks, it demonstrated a more effective exploration strategy, discovering a crucial 10% more of the critical failure region within the same budget. Most notably, in tasks with unknown user preferences, our approach always successfully learned the underlying objective where other baselines fail to capture. Together, these findings validate the power of our unified approach, showing that a principled balance between pragmatic and epistemic drives leads to tangible gains across diverse and challenging problem settings.
In summary, our main contributions are the following:
  • A unified perspective of various acquisition strategies through the lens of active inference.
  • A pragmatic curiosity paradigm for general hybrid learning-optimization problems.
  • A comprehensive empirical validation across three canonical real-world problem categories with diverse hybrid learning-optimization objectives.
@@ 2.1 Bayesian Optimization
Given an unknown objective function $f:\mathcal{X}\mapsto\Re$, BO seeks to identify the input $x^{\ast}$ that maximizes the objective $f$ over an admissible set of queries $\mathcal{X}$, i.e., $x^{\ast}=\arg\max_{x\in\mathcal{X}}f(x).$ To achieve this goal, BO relies on a surrogate model that provides a probabilistic representation of the objective $f$, and uses this information to compute an acquisition function to drive the selection of the most promising sample to query.
Surrogate model. We assume the available information regarding the objective function $f$ be stored in the dataset $\mathcal{D}_{t}:=\{(x_{1},y_{1}),\dots,(x_{t},y_{t}))\}$, where $y_{t}\sim\mathcal{N}(f(x_{t}),\sigma^{2}(x_{t}))$ is the noisy observation of the objective function by assuming the noise follows a zero-mean normal distribution with a standard deviation $\sigma$. The surrogate model depicts possible explanations of $f$ as $f(x)\sim p(f(x)|\mathcal{D}_{t})$ applying a joint distribution over its behavior at each sample $x\in\mathcal{X}$. In Bayesian inference, the prior distribution of the objective $p(f(x))$ is combined with the likelihood function $p(\mathcal{D}_{t}|f(x))$ to compute the posterior distribution $p(f(x)|\mathcal{D}_{t})\propto p(\mathcal{D}_{t}|f(x))p(f(x))$, representing the updated beliefs about $f(x)$. Typically, Gaussian processes (GPs) have been widely used as the surrogate model for BO due to their efficient posterior sampling that enables cheap, gradient-based optimization of the acquisition function to propose new query points. GP is specified by a joint normal distribution $p(f(x)|\mathcal{D}_{t})=\mathcal{N}(\mu_{t}(x),\kappa_{t}(x,x^{\prime}))$ with mean $\mu_{t}(x)$ and kernel function $\kappa_{t}(x,x^{\prime})$, where $\mu_{t}(x)$ represents the prediction and $\kappa_{t}(x,x^{\prime})$ the associated uncertainty.
Acquisition function. The surrogate model is utilized to decide the next sample $x_{t+1}\in\mathcal{X}$ through the maximization of an acquisition function $\alpha:\mathcal{X}\mapsto\Re$, i.e., $x_{t+1}=\arg\max_{x\in\mathcal{X}}\alpha(x|\mathcal{D}_{t})$, where $\alpha(x|\mathcal{D}_{t})$ provides a measure of the improvement that the next query is likely to provide with respect to (w.r.t.) the current surrogate model of the objective function. Many acquisition functions have been proposed, including probability of improvement (Močkus, 1975), expected improvement (Jones et al., 1998), upper confidence bound, and various entropy search methods (Hennig and Christian J. Schuler, 2012; Hernández-Lobato et al., 2014; Wang and Jegelka, 2017; Hvarfner et al., 2022b; Neiswanger et al., 2021), as well as practical approaches to optimize them (Wilson et al., 2018).
@@ 2.2 Bayesian Experimental Design
Based on the current history of experiments $\mathcal{D}_{t}:=\{(x_{1},y_{1}),\dots,(x_{t},y_{t}))\}$, BED seeks to find the next experimental design $x_{t+1}$ by maximizing the expected information gain (EIG) (Chaloner and Verdinelli, 1995) that a potential experimental outcome $y_{t+1}$ can provide about $\theta$, measured in terms of expected entropy reduction of the posterior distribution of $\theta$:
$\begin{split}\text{EIG}(x|\mathcal{D}_{t})&=H[p(\theta|\mathcal{D}_{t})]-\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[H[p(\theta|\mathcal{D}_{t}\cup(x,y))]]\\ &=I(\theta;(x,y)|\mathcal{D}_{t}),\end{split}$
@@
3 A Unified View of Acquisition Strategies
The acquisition strategies in BO typically lead to goal-directed behavior, where the (implicit) goal is the optimum of a certain (unknown) objective function. On the contrary, the acquisition strategies in BED encourage information-seeking behavior, aiming to gather the maximum amount of information about certain parameters of interest. Although they both can be viewed as realizations of adaptive sampling (Di Fiore et al., 2023), there are no transferable approaches between these two domains due to distinct directives (Hvarfner et al., 2025).
In this section, we show that these two seemingly competing imperatives can be naturally balanced through the principles of active inference (AIF).
@@ 3.1 Active Inference as Expected Free Energy Minimization
We specify a probabilistic surrogate model $q(\cdot)$ to capture the relationship between an outcome $y$ and a decision variable $x$ based on a set of parameters $s$ that are of interest, which factorizes as
$q(x,y,s)\coloneq p(x,y|s)q(s),$
where $q(x,y,s)$ is a joint probability distribution over $(x,y,s)$ based on a surrogate distribution $q(s)$ over $s$. We use $q(\cdot)$ to explicitly distinguish this model to be an “surrogate” of the “true” model $p(\cdot)$.
One special aspect of AIF is the way it formalizes goals. Instead of specifying the goal with additional variables related to “rewards” or “costs”, AIF directly encodes the preferences over possible outcomes $y$ through a probability distribution $p(y)$. In this distribution, outcomes with higher probabilities are treated as more rewarding. The deviation between observed outcomes and those desired is measured through an information-theoretic quantity known as self-information or surprisal: $\log p(y)$. Intuitively, it’s a measure of how unexpected an outcome $y$ is, given a prior preference distribution $p(y)$. Consistent with the intuitive notion of surprise, lower probability events generate higher surprisal values. Then the surprisal incurred from an observed outcome $y$ is quantified by
$\begin{split}-\log p(y)&=-\log\int p(y,s)ds\\ &=-\log\int\frac{p(y,s)q(s)}{q(s)}ds\\ &=-\log\mathbb{E}_{q(s)}\left[\frac{p(y,s)}{q(s)}\right]\\ &\leq-\mathbb{E}_{q(s)}\left[\log\frac{p(y,s)}{q(s)}\right]=F,\end{split}$ (1)
where the last inequality follows Jensen’s inequality, which states that the expectation of a logarithm is always less than or equal to the logarithm of an expectation.
@@
The right-hand side of (1) is called the variational free energy (VFE), whose name arises from the fact that $F$ resembles the Helmholtz free energy in physics. We can see that VFE is always greater than or equal to surprisal (i.e., it is an upper bound on surprisal). In machine learning, the sign of VFE is typically reversed, making it an evidence lower bound (ELBO). Maximizing the ELBO is a commonly used optimization approach in machine learning (Titsias, 2009).
To formulate a strategy for decision-making, we need to consider the decision variable $x$ and the outcomes that result from the choice of actions. Since the future outcomes have not yet occurred, we resort to examining the expectation of surprisal over predicted outcomes based on a predictive distribution $q(y|x)$:
$\begin{split}-\mathbb{E}_{q(y|x)}&\log p(y|x)\\ \leq-&\mathbb{E}_{q(y,s|x)}\left[\log\frac{p(y,s|x)}{q(s|x)}\right]\\ =-&\mathbb{E}_{q(y,s|x)}\left[\log\frac{p(s|x,y)p(y|x)}{q(s|x)}\right]\\ =-&\mathbb{E}_{q(y,s|x)}\left[\log p(s|x,y)-\log q(s|x)\right]\\ &-\mathbb{E}_{q(y,s|x)}\log p(y|x)\\ =-&\mathbb{E}_{q(y,s|x)}\left[\log p(s|x,y)-\log q(s|x)\right]\\ &-\mathbb{E}_{q(y|x)}\log p(y|x)=G,\end{split}$ (2)
@@
where the right-hand side of (2) is denoted as the expected free energy (EFE).
Theorem 3.1.
When using a surrogate model $q(\cdot)=p(\cdot|\mathcal{D}_{t})$ that is considered as the true model constructed from all available data $\mathcal{D}_{t}$, $G$ can be simplified as

### v2

# Pragmatic Curiosity: A Unified Framework for Hybrid Learning and Optimization via Active Inference

@@ Abstract
Many engineering and scientific workflows rely on expensive black-box evaluations, requiring sequential decisions that must both improve task performance and reduce uncertainty. Bayesian optimization (BO) and Bayesian experimental design (BED) provide powerful but largely separate treatments of goal-directed optimization and information-seeking experimentation, leaving limited guidance for hybrid settings in which learning and optimization are intrinsically coupled. We propose Pragmatic Curiosity (PraC), a unified framework for hybrid learning and optimization via active inference. PraC evaluates candidate queries by trading information gain about a task-relevant latent symbol against an expected regret-based potential over outcomes. This formulation exposes three operational design choices: which latent quantity should be clarified, how task value is encoded as regret, and how strongly information gain should be exchanged against pragmatic value. We instantiate PraC across three regimes of increasing complexity: decision-oriented plume monitoring with fixed global symbols and known downstream losses, targeted active search with induced local symbols and evolving coverage goals, and composite Bayesian optimization with hierarchical regret learning under unknown preferences. Across these regimes, PraC reduces downstream decision risk, improves coverage of critical outcome regions, and jointly learns predictive and preference structures without relying on task-specific staging rules.
@@ 1 Introduction
Sequential decision-making under uncertainty often requires an agent to decide what to learn and what to do at the same time. Two classical paradigms have addressed these aspects from different directions. Bayesian optimization (BO) is primarily goal-directed: it selects queries in order to identify high-performing inputs of an unknown objective function (Močkus, 1975b; Jones et al., 1998a; Srinivas et al., 2012). Bayesian experimental design (BED) is primarily information-seeking: it selects experiments in order to reduce uncertainty about latent quantities of interest (Chaloner and Verdinelli, 1995a). Although both can be viewed as forms of adaptive sampling (Di Fiore et al., 2023), they are usually developed under different directives, with different acquisition rules and different criteria for success.
This separation leaves limited guidance for problems in which the relevant uncertainty, the downstream objective, and the sampling policy must be designed jointly. In environmental monitoring (Konakovic Lukovic et al., 2020), a robot may need to learn the hidden state of a chemical plume while deciding where to dispatch a response team. In failure discovery (Ramanagopal et al., 2018; Parashar et al., 2025), an evaluator may need to explore uncertain scenarios while prioritizing outcomes that cover critical failure regions. In preference-guided design (González and Zavala, 2025; Coelho et al., 2025), an optimizer may need to learn not only how actions produce outcomes, but also how those outcomes should be valued. In all these settings, an agent must act before it fully knows, and it only learns by acting. The bottleneck is therefore not learning or optimization alone, but the meta-decision that connects them: deciding what knowledge is sufficient for action, and what actions are most informative for future decisions.
@@
Recent work has begun to address this coupling by choosing between specialized tools and accommodating problem-specific adaptation through leveraging information-gain criteria to enhance optimization and vice versa. Information-directed sampling (IDS) connects information and regret in online optimization and bandit settings (Russo and Van Roy, 2018). Self-correcting Bayesian optimization introduces a statistical distance-based active learning criterion into the BO loop to improve model learning during optimization (Hvarfner et al., 2023). Expected predictive information gain (EPIG) focuses active learning on predictive information that is relevant to a target distribution (Smith et al., 2023). Preference exploration for BO learns preference models over multi-objective outcomes, but typically relies on stage-wise choices between experimentation and preference learning (Jerry Lin et al., 2022). These methods highlight growing synergy between learning and optimization, but they often remain task-specific, and rarely generalize across categories.
In this paper, we propose Pragmatic Curiosity (PraC), a unified framework for hybrid learning and optimization via active inference (AIF) (Friston, 2010; Friston et al., 2017). Originally developed in computational neuroscience, AIF prescribes action selection by minimizing expected free energy (EFE), a single objective that couples an epistemic drive for information gain with an pragmatic drive toward preferred outcomes. PraC translates this principle into a practical acquisition-design template: a candidate query is evaluated by both what it is expected to reveal about a task-relevant latent quantity and how its predicted outcome is expected to affect downstream action. In this sense, PraC formalizes pragmatic curiosity: curiosity is not the indiscriminate gathering of information, but the search for knowledge insofar as that knowledge can improve action.
PraC not only serves as a unifying umbrella principle that re-interprets many classical acquisition rules in BO and BED, but also opens the door to practical acquisition design for a broader class of hybrid problems by exposing a triad of meta-decisions in such problems. First, the agent must decide what hidden structure is worth representing, since not all uncertainty is relevant to action. Second, it must decide how outcomes should be evaluated, since different downstream goals induce different notions of success, failure, risk, or shortfall. Third, it must decide how strongly unresolved uncertainty should influence the next action, since acting before knowing can be brittle while learning without regard to the goal can become aimless. Thus, hybrid decision-making is governed by a triad of meta-decisions in representation, evaluation, and regulation.
In PraC, this triad is conceptualized as symbols, regrets, and curiosity. Symbols specify the latent quantities whose clarification can change what the agent ought to do. In parametric models, these may be fixed global symbols such as physical parameters or latent hypotheses; in non-parametric models, they may be induced local symbols tied to contemplated observations. Regret acts as a carrier of goal: it gives an operational representation of what matters by quantifying how an outcome falls short of what is desired. When the goal is known, a regret-based potential can be designed directly to encode domain knowledge; when the goal is unknown or implicit, the regret itself can be learned through a hierarchical structure of PraC. Curiosity specifies the exchange rate between clarifying knowledge and pursuing the current surrogate of the goal. Together, these three choices turn PraC from an abstract decision paradigm into a practical decision rule that accounts for diverse hybrid learning and optimization problems.
This perspective also elevates curiosity from an ad-hoc exploration heuristic into an intrinsic regularizer that couples belief updating and decision-making. In many acquisition functions, the exploration weight appears as a fixed hyperparameter selected by empirical tuning. In PraC, curiosity has a semantic role: it determines how much pragmatic sacrifice is justified by epistemic clarification. This allows the effective epistemic pressure of the acquisition to be monitored and scheduled dynamically. We introduce a practical feedforward–feedback scheduler that calibrates curiosity from the relative scale of information gain and downstream pragmatic value, while suppressing curiosity as decision-relevant uncertainty decreases. In this way, curiosity becomes a fluid exchange rate that self-regulates as belief and action co-evolve.
We instantiate PraC across three regimes of increasing complexity and empirically evaluate the role of each design axis. First, we study fixed global symbols with known downstream goals in decision-oriented monitoring. Here the latent symbol is a finite hypothesis, and the regret is the Bayes risk of a downstream response decision. PraC with dynamic curiosity scheduler achieves the lowest final Bayes risk in source response localization and active source prioritization, and obtains the best realized response loss in consequence-weighted dispatch, showing that curiosity can help resolve symbolic distinctions that matter for action rather than merely reducing uncertainty for its own sake.
Second, we study induced local symbols with known evolving goals in targeted active search. Here the relevant symbol is induced by a non-parametric model, and the pragmatic value of a new outcome depends on what has already been covered. PraC coordinates exploration in the input space with coverage in the outcome space, and in the most challenging target set discovers nearly 10% more normalized target-region coverage than the strongest baseline (Malkomes et al., 2021a). This demonstrates the benefit of allowing symbols to be induced locally by contemplated observations rather than fixed globally in advance.
Third, we study hierarchical symbols with unknown goals in composite Bayesian optimization with preferences. Here the agent must learn both how actions produce outcomes and how outcomes should be valued. PraC consistently outperforms ablated variants that remove components of the hierarchical acquisition, and further improves over BOPE-style stage-wise preference-exploration variants (Jerry Lin et al., 2022). These results show that a hierarchical regret design can jointly support outcome learning and preference learning without manually separating them into stages.
Our contributions are as follows.
  • We derive a practical active-inference acquisition template that couples information gain about a task-relevant latent symbol with an expected regret-based potential over outcomes.
  • We identify three operational design axes—epistemic symbol, regret potential, and curiosity coefficient—and show how they recover, reinterpret, or approximate several BO/BED acquisition strategies while supporting new hybrid designs.
  • We introduce a dynamic curiosity scheduler that calibrates the information–regret exchange rate from candidate-level acquisition statistics and posterior uncertainty.
  • We demonstrate the framework across decision-oriented environmental monitoring, targeted active search, and preference-guided composite optimization, showing that the same principle adapts across parametric, non-parametric, and hierarchical regimes.
@@ 2.1 Bayesian Optimization
Given an unknown objective function $f:\mathcal{X}\to\mathbb{R}$, BO seeks to identify the input $x^{\ast}$ that maximizes the objective $f$ over an admissible set of queries $\mathcal{X}$, i.e., $x^{\ast}=\arg\max_{x\in\mathcal{X}}f(x).$ To achieve this goal, BO relies on a surrogate model that provides a probabilistic representation of the objective $f$, and uses this information to compute an acquisition function to drive the selection of the most promising sample to query.
Surrogate model. We assume the available information regarding the objective function $f$ be stored in the dataset $\mathcal{D}_{t}:=\{(x_{1},y_{1}),\dots,(x_{t},y_{t})\}$, where $y_{t}\sim\mathcal{N}(f(x_{t}),\sigma^{2}(x_{t}))$ is the noisy observation of the objective function by assuming the noise follows a zero-mean normal distribution with a standard deviation $\sigma(x)$. The surrogate model depicts possible explanations of $f$ as $f(x)\sim p(f(x)|\mathcal{D}_{t})$ applying a joint distribution over its behavior at each sample $x\in\mathcal{X}$. In Bayesian inference, the prior distribution of the objective $p(f(x))$ is combined with the likelihood function $p(\mathcal{D}_{t}|f(x))$ to compute the posterior distribution $p(f(x)|\mathcal{D}_{t})\propto p(\mathcal{D}_{t}|f(x))p(f(x))$, representing the updated beliefs about $f(x)$. Typically, Gaussian processes (GPs) have been widely used as the surrogate model for BO due to their efficient posterior sampling that enables cheap, gradient-based optimization of the acquisition function to propose new query points. GP is specified by a joint normal distribution $p(f(x)|\mathcal{D}_{t})=\mathcal{N}(\mu_{t}(x),\kappa_{t}(x,x^{\prime}))$ with mean $\mu_{t}(x)$ and kernel function $\kappa_{t}(x,x^{\prime})$, where $\mu_{t}(x)$ represents the prediction and $\kappa_{t}(x,x^{\prime})$ the associated uncertainty.
Acquisition function. The surrogate model is utilized to decide the next sample $x_{t+1}\in\mathcal{X}$ through the maximization of an acquisition function $\alpha:\mathcal{X}\to\mathbb{R}$, i.e., $x_{t+1}=\arg\max_{x\in\mathcal{X}}\alpha(x\mid\mathcal{D}_{t})$, where $\alpha(x\mid\mathcal{D}_{t})$ provides a measure of the improvement that the next query is likely to provide with respect to (w.r.t.) the current surrogate model of the objective function. Many acquisition functions have been proposed, including probability of improvement (Močkus, 1975a), expected improvement (Jones et al., 1998b), upper confidence bound, and various entropy search methods (Hennig and Christian J. Schuler, 2012; Hernández-Lobato et al., 2014; Wang and Jegelka, 2017; Hvarfner et al., 2022; Neiswanger et al., 2021), as well as practical approaches to optimize them (Wilson et al., 2018).
@@ 2.2 Bayesian Experimental Design
Based on the current history of experiments $\mathcal{D}_{t}:=\{(x_{1},y_{1}),\dots,(x_{t},y_{t})\}$, BED seeks to find the next experimental design $x_{t+1}$ by maximizing the expected information gain (EIG) (Chaloner and Verdinelli, 1995b) that a potential experimental outcome $y_{t+1}$ can provide about $\theta$, measured in terms of expected entropy reduction of the posterior distribution of $\theta$:
$\begin{split}\text{EIG}(x\mid\mathcal{D}_{t})&=H[p(\theta|\mathcal{D}_{t})]-\mathbb{E}_{p(y\mid x,\mathcal{D}_{t})}[H[p(\theta|\mathcal{D}_{t}\cup(x,y))]]=I(\theta;(x,y)\mid\mathcal{D}_{t}),\end{split}$
@@
3 Pragmatic Curiosity: A Framework that Synthesizes Learning and Optimization
Acquisition strategies in BO typically induce goal-directed behavior, where the implicit goal is to identify the optimum of an unknown objective function. By contrast, acquisition strategies in BED induce information-seeking behavior, aiming to collect the most informative data about latent quantities of interest. Although both can be viewed as instances of adaptive sampling (Di Fiore et al., 2023), they are usually developed under different directives and therefore admit few directly transferable design principles (Hvarfner et al., 2025). In this section, we synthesize these two seemingly competing imperatives within a unified framework through the lens of active inference (AIF).
@@ 3.1 Active Inference as Expected Free Energy Minimization
Consider a probabilistic surrogate model $q(\cdot)$ that captures the relationship between a decision variable $x$, an outcome $y$, and a collection of latent quantities of interest $s$, and factorizes as
$q(x,y,s)\coloneq p(x,y\mid s)\,q(s),$
where $q(s)$ is the surrogate belief over $s$. We use $q(\cdot)$ to distinguish the agent’s internal surrogate model from the underlying, generally inaccessible, true data-generating model $p(\cdot)$.
In AIF, preferences over possible outcomes are encoded through a probability distribution $p(y)$. Outcomes assigned higher probability are treated as more preferred. The deviation between an observed outcome and those preferred by the agent is measured by its self-information, or surprisal, $\log p(y)$. Intuitively, surprisal quantifies how unexpected an outcome $y$ is under the preference distribution $p(y)$: less preferred outcomes incur larger surprisal.
The surprisal associated with an outcome $y$ satisfies
$\begin{split}-\log p(y)&=-\log\int p(y,s)\,ds=-\log\int\frac{p(y,s)q(s)}{q(s)}\,ds\\ &=-\log\mathbb{E}_{q(s)}\!\left[\frac{p(y,s)}{q(s)}\right]\leq-\mathbb{E}_{q(s)}\!\left[\log\frac{p(y,s)}{q(s)}\right]=F,\end{split}$ (1)
where the inequality follows from Jensen’s inequality.
@@
The right-hand side of equation 1 is called the variational free energy (VFE), resembling the Helmholtz free energy in physics. It upper bounds surprisal and therefore provides a tractable surrogate objective for inference. In machine learning, the sign of VFE is often reversed, yielding the evidence lower bound (ELBO), whose maximization is a standard variational learning principle (Titsias, 2009).
To obtain a decision rule, we must account not only for realized outcomes, but also for future outcomes that have not yet been observed. AIF does so by considering the expected surprisal of future outcomes under the predictive distribution $q(y\mid x)$:
$\begin{split}-\mathbb{E}_{q(y\mid x)}&\log p(y\mid x)\leq-\mathbb{E}_{q(y,s\mid x)}\left[\log\frac{p(y,s\mid x)}{q(s\mid x)}\right]=-\mathbb{E}_{q(y,s\mid x)}\left[\log\frac{p(s\mid x,y)p(y\mid x)}{q(s\mid x)}\right]\\ =-&\mathbb{E}_{q(y,s\mid x)}\left[\log p(s\mid x,y)-\log q(s\mid x)\right]-\mathbb{E}_{q(y,s\mid x)}\log p(y\mid x)\\ =-&\mathbb{E}_{q(y,s\mid x)}\left[\log p(s\mid x,y)-\log q(s\mid x)\right]-\mathbb{E}_{q(y\mid x)}\log p(y\mid x)=G,\end{split}$ (2)
@@
where the right-hand side of equation 2 is denoted as the expected free energy (EFE). AIF prescribes action selection by minimizing this quantity.
3.2 A Principled Decision Paradigm for Hybrid Learning and Optimization
In its original form, the EFE in equation 2 is not yet an actionable acquisition rule, because it depends on the true posterior $p(s\mid x,y)$ and the true outcome model $p(y\mid x)$, both of which are generally unavailable a priori. We now turn it into a principled decision paradigm through two approximations that are natural from a decision-theoretic perspective.
The EFE in equation 2 can be decomposed as
$G=\underbrace{-\mathbb{E}_{q(y,s\mid x)}\left[\log p(s\mid x,y)-\log q(s\mid x)\right]}_{\text{Term 1}}\;\underbrace{-\mathbb{E}_{q(y\mid x)}\log p(y\mid x)}_{\text{Term 2}}.$
For Term 1, add and subtract $\log q(s\mid x,y)$ inside the expectation: